# revtex4-1Repair the float

### v1

@@ IV Conclusions
In conclusion, we have considered few-site Kitaev chains and investigated the emergence of superconducting pair correlations. Under general circumstances, we have shown that distinct pair symmetries are allowed to form in these systems as a result of the multiple existing quantum numbers. In the case of a two-site Kitaev chain, we found local pairing with an odd-frequency dependence, while nonlocal pair correlations with both even- and odd-frequency components. While in general the odd-frequency components are linear, we discovered that they acquire a divergent profile around zero frequency when the nonlocal pair potential and electron tunneling are of the same order and at least one onsite energy tuned to zero. Since these regime corresponds to the phase with poor man’s Majorana modes, we have interpreted the divergent odd-frequency pairing as a signature of charge neutrality and spatial nonlocality, intrinsic to Majorana quasiparticles. The results presented here can be of use to understand the nature of superconducting correlations in few-site Kitaev chains.
@@ V Acknowledgements
We thank R. Seoane Souto for insightful discussions. We acknowledge financial support from the Swedish Research Council (Vetenskapsrådet Grant No. 2021-04121), the Carl Trygger’s Foundation (Grant No. 22: 2093), and the Göran Gustafsson Foundation (Grant No. 2216).

### v2

@@ IV Conclusions
In conclusion, we have considered few-site Kitaev chains and investigated the emergence of superconducting pair correlations. Under general circumstances, we have shown that distinct pair symmetries are allowed to form in these systems as a result of the multiple existing quantum numbers. In the case of a two-site Kitaev chain, we found local pairing with an odd-frequency dependence, while nonlocal pair correlations with both even- and odd-frequency components. While in general the odd-frequency components are linear, we discovered that they acquire a divergent profile around zero frequency in a fine-tuned sweet spot occurring when the parent pair potential and electron tunneling are of the same order and at least one onsite energy tuned to zero. Since these sweet spot regime corresponds to the phase with poor man’s Majorana modes, which are fully nonlocal, self-conjugate, and charge neutral, we have interpreted the divergent odd-frequency pairing as a signature of charge neutrality and spatial nonlocality, intrinsic to Majorana quasiparticles. The results presented here can be of use to understand the nature of superconducting correlations in few-site Kitaev chains.
@@ V Acknowledgements
We thank R. Seoane Souto for insightful discussions. We acknowledge financial support from the Swedish Research Council (Vetenskapsrådet Grant No. 2021-04121), the Carl Trygger’s Foundation (Grant No. 22: 2093), and the Göran Gustafsson Foundation (Grant No. 2216). The computations were enabled by resources provided by the National Academic Infrastructure for Supercomputing in Sweden (NAISS), partially funded by the Swedish Research Council through Grant Agreement No. 2022-06725.